# Heavy-Tailed Linear Bandits: Huber Regression with One-Pass Update

### v1

@@ 5.1 Online Linear MDPs
In the setting of heavy-tailed linear MDPs (Li and Sun, 2023; Huang et al., 2024), the realizable reward for taking action $a$ in state $s$ at the $h$-th episode is given by:
@@
Linear MDPs often adopt techniques from linear bandits, including estimating unknown parameters in a linear model and constructing an upper confidence bound (UCB) based on estimation error analysis. For heavy-tailed linear MDPs, (Huang et al., 2024) employs adaptive Huber regression to estimate the unknown parameter $\bm{\theta}_{h}^{*}$ in the reward function. This introduces a challenge similar to that in HvtLB: adaptive Huber regression is an offline algorithm, while an online setting requires a one-pass estimator. Existing one-pass estimators for HvtLB, such as the truncation-based and MOM-based algorithms proposed by Xue et al. (2023), are not applicable to heavy-tailed linear MDPs. Truncation-based algorithms rely on assumptions about absolute moments, which are not suitable in this context. MOM-based methods require multiple reward observations for a fixed $\phi(s,a)$, but this is infeasible in linear MDPs since state transitions occur probabilistically, preventing repeated observations for the same state-action pair.
In contrast, our proposed Hvt-UCB method can be directly applied to this scenario without requiring additional assumptions. Its estimation error bound analysis naturally extends to linear MDPs, enabling the construction of UCBs for the value function. By leveraging the UCB-to-regret analysis in (Huang et al., 2024), Hvt-UCB has the potential to achieve the same theoretical guarantees as adaptive Huber regression, while significantly reducing computational costs.
@@ 5.2 Online Adaptive Control
The online adaptive control of Linear Quadratic Systems (Abbasi-Yadkori and Szepesvári, 2011) considers the following state transition system,
@@
where $x_{t}\in\mathbb{R}^{n}$ represents the state at time $t$, $u_{t}\in\mathbb{R}^{d}$ is the control input at time $t$, $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times d}$ are unknown system matrices, and $w_{t+1}$ is the noise. The online adaptive control requires estimating the unknown parameters $A$ and $B$ (also known as system identification), and constructing finite-sample guarantees for estimation error. To this end, Abbasi-Yadkori and Szepesvári (2011) transformed the system into the following linear model:
@@
As noted by Tsiamis et al. (2023), finite-sample guarantees for system identification under heavy-tailed noise remain an open challenge. Inspired by Abbasi-Yadkori and Szepesvári (2011)’s application of linear bandit techniques to sub-Gaussian cases, we believe methods from HvtLB can similarly benefit adaptive control in heavy-tailed settings. However, MOM methods are infeasible as the feature $z_{t}$, which includes the evolving state $x_{t}$, cannot be repeatedly sampled, and truncation-based methods rely on assumptions of bounded absolute moments for states, which may not hold. While adaptive Huber regression (Li and Sun, 2023; Huang et al., 2024) offers promising theoretical guarantees for heavy-tailed linear system identification, its computational inefficiency makes it unsuitable for adaptive control. In contrast, our proposed Hvt-UCB algorithm provides efficient, one-pass updates without additional assumptions and has the potential to deliver finite-sample guarantees for adaptive control under heavy-tailed noise.
6 Experiments
In this section, we empirically evaluate the performance and time efficiency of our proposed algorithm under two different noise distributions: Student’s $t$ and Gaussian.
Settings.   We consider the linear model $r_{t}=X_{t}^{\top}\theta_{*}+\eta_{t}$, where the dimension $d=2$, the number of rounds $T=18000$, and the number of arms $n=50$. Each dimension of the feature vectors for the arms is uniformly sampled from $[-1,1]$ and subsequently rescaled to satisfy $L=1$. Similarly, $\theta_{*}$ is sampled in the same way and rescaled to satisfy $S=1$. We conduct two synthetic experiments with different distributions for noise $\eta_{t}$: (a) Student’s $t$-distribution with degree of freedom $df=2.1$ to represent heavy-tailed noise; and (b) Gaussian noise sampled from $\mathcal{N}(0,1)$ to represent light-tailed noise. For the heavy-tailed experiment, we set $\varepsilon=0.99$ and $\nu=1.31$, while for the light-tailed experiment, we set $\varepsilon=1$. We compare the performance of our proposed Hvt-UCB algorithm with the following baselines: (a) the OFUL algorithm (Abbasi-Yadkori et al., 2011); (b) the one-pass truncation-based algorithm CRTM (Xue et al., 2023); (c) the one-pass MOM-based algorithm CRMM (Xue et al., 2023); and (d) the Huber-based algorithm HEAVY-OFUL (Huang et al., 2024).
Figure 1: Student $t$: regret and time cost.
Figure 2: Gaussian: regret and time cost.
Results.   We conducted 10 independent trials and averaged the results. Figure 1 shows the cumulative regret and computational time under Student’s $t$ noise, while Figure 2 presents the same metrics under Gaussian noise, with shaded regions indicating the variance across trials. Under both noise settings, our algorithm demonstrates comparable regret performance to Heavy-OFUL while achieving remarkable computational efficiency, with a speedup factor exceeding 800×800$\times$800 ×. This efficiency advantage becomes more important as the number of rounds $T$ increases. Notably, among the one-pass algorithms, CRMM exhibits the fastest runtime because it updates only every $\mathcal{O}(\log T)$ rounds rather than at each round. In the heavy-tailed noise scenario (Figure 1), the performance of OFUL drops significantly, demonstrating its sensitivity to heavy-tailed distributions. In the Gaussian noise setting (Figure 2), our algorithm maintains competitive performance alongside OFUL. However, the truncation-based CRTM algorithm shows suboptimal performance due to excessive reward truncation in light-tailed noise scenarios, which leads to the loss of valuable data. These experimental results validate that our Hvt-UCB algorithm achieves robust performance across different noise environments while maintaining low computational cost.
@@ 7 Conclusion
In this paper, we investigate the heavy-tailed linear bandits problem. We identify the advantages of Huber loss-based methods over the truncation and median-of-means strategies, while also pointing out the inefficiencies of previous Huber loss-based methods. We then propose the Huber loss-based one-pass algorithm Hvt-UCB based on the Online Mirror Descent (OMD) framework. Hvt-UCB achieves the optimal and instance-dependent regret bound $\widetilde{\mathcal{O}}\big{(}dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\sqrt{\sum_{t=1}^{T}\nu_{t}^{2}}+dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\big{)}$ while only requiring $\mathcal{O}(1)$ per-round computational cost. Furthermore, it can be extended to online linear MDPs and online adaptive control, broadening its applicability. The key contribution is our adaptation of the OMD framework to stochastic linear bandits, which addresses the challenges introduced by heavy-tailed noise and the structure of the Huber loss.
Both Huang et al. (2024) and our work rely on knowledge of the moment $\nu_{t}$ to achieve instance-dependent regret guarantees. Extending these Huber loss-based methods to handle unknown $\nu_{t}$ while maintaining instance-dependent guarantees is an interesting problem for future investigation.
@@ Appendix A Properties of Huber Loss
Property 1 (Property 1 of Huang et al. [2024]).
@@ B.2 Proof of Lemma 3
$\begin{split}{}&\frac{1}{2}\left\|\widehat{\theta}_{t}-\theta_{*}\right\|^{2}_{\frac{X_{t}X_{t}^{\top}}{\sigma_{t}^{2}}}\\ \leq{}&\left\langle\nabla\ell_{t}(\widehat{\theta}_{t}),\widehat{\theta}_{t+1}-\theta_{*}\right\rangle+\left\langle\nabla\ell_{t}(\widehat{\theta}_{t}),\widehat{\theta}_{t}-\widehat{\theta}_{t+1}\right\rangle+\left\langle\nabla\widetilde{\ell}_{t}(\widehat{\theta}_{t})-\nabla\ell_{t}(\widehat{\theta}_{t}),\widehat{\theta}_{t}-\theta_{*}\right\rangle,\end{split}$ (21)
where the first term can be bounded by the Bregman proximal inequality (9), we have
@@ B.3 Proof of Lemma 4
By appling Lemma C.5 of Huang et al. [2024], restated in Lemma 6, with $\tau_{t}=\tau_{0}\frac{\sqrt{1+w_{t}^{2}}}{w_{t}}t^{\frac{1-\varepsilon}{2(1+\varepsilon)}}$ and $b=\max_{t\in[T]}\frac{\nu_{t}}{\sigma_{t}}\leq 1$, we have with probability at least $1-\delta,\forall t\geq 1$,
@@ B.4 Proof of Lemma 5
where the last inequality comes from that $\max_{s\in[t]}\beta_{s-1}\leq\max_{s\in[t+1]}\beta_{s-1}$ Same as Eq. (C.12) of Huang et al. [2024], by setting $\tau_{0}$ as (17), with probability at least $1-\delta$, for all $t\geq 1$, we have
@@
where the last inequality comes from that $\operatorname{\mathds{1}}_{A_{s}}\leq 1$. Then, Lemma C.2 of Huang et al. [2024] (Self-normalized concentration) shows that by setting $\tau_{0}$ as (17) and $b=\max_{t\in[T]}\frac{\nu_{t}}{\sigma_{t}}\leq 1$, we have with probability at least $1-2\delta$, $\forall t\geq 1$,
@@ C.1 Proof of Theorem 1
where the last inequality comes from Lemma 7. Combine these two cases together, by choosing $\delta=\frac{1}{8T}$, $\lambda=d$, we have
@@ Appendix D Technical Lemmas
This section contains some useful technical lemmas that are used in the proofs.
@@ D.1 Concentrations
Theorem 2 (Self-normalized concentration for scalar [Abbasi-Yadkori et al., 2012, Lemma 7]).
@@
Lemma 6 (Lemma C.5 of Huang et al. [2024]).
@@ D.2 Potential Lemma
Note that in the last inequality, we utilize the fact that $\lVert U_{t-1}^{-\frac{1}{2}}X_{t}\rVert_{2}^{2}\leq c_{\max}$ and $1+x\geq\exp(x/2)$ holds for any $x\in[0,1]$. By taking advantage of the telescope structure, we have
@@
Lemma 8 (Lemma 5 of Zhao et al. [2020]).
@@ D.3 Useful Lemma for OMD
Lemma 9 (Bregman proximal inequality [Chen and Teboulle, 1993, Lemma 3.2]).

### v2

@@ 6.1 Online Linear MDPs
In the setting of heavy-tailed linear MDPs (Li & Sun, 2024; Huang et al., 2023), the realizable reward for taking action $a$ in state $s$ at the $h$-th episode is given by:
@@
Linear MDPs often adopt techniques from linear bandits, including estimating unknown parameters in a linear model and constructing an upper confidence bound (UCB) based on estimation error analysis. For heavy-tailed linear MDPs, (Huang et al., 2023) employs adaptive Huber regression to estimate the unknown parameter $\bm{\theta}_{h}^{*}$ in the reward function. This introduces a challenge similar to that in HvtLB: adaptive Huber regression is an offline algorithm, while an online setting requires a one-pass estimator. Existing one-pass estimators for HvtLB, such as the truncation-based and MOM-based algorithms proposed by Xue et al. (2023), are not applicable to heavy-tailed linear MDPs. Truncation-based algorithms rely on assumptions about absolute moments, which are not suitable in this case. MOM-based methods require repeated reward observations for a fixed $\phi(s,a)$, which is infeasible in linear MDPs since state transitions occur probabilistically.
In contrast, our proposed Hvt-UCB method does not suffer the aforementioned issue and has the potential to be applied to this scenario without requiring additional assumptions. Its estimation error bound analysis naturally extends to linear MDPs, enabling the construction of UCBs for the value function. By leveraging the UCB-to-regret analysis in (Huang et al., 2023), Hvt-UCB has the potential to achieve the same theoretical guarantees as adaptive Huber regression, while significantly reducing computational costs.
@@ 6.2 Online Adaptive Control
The online adaptive control of Linear Quadratic Systems (Abbasi-Yadkori & Szepesvári, 2011) considers the following state transition system,
@@
where $x_{t}\in\mathbb{R}^{n}$ represents the state at time $t$, $u_{t}\in\mathbb{R}^{d}$ is the control input at time $t$, $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times d}$ are unknown system matrices, and $w_{t+1}$ is the noise. The online adaptive control requires estimating the unknown parameters $A$ and $B$ (also known as system identification), and constructing finite-sample guarantees for estimation error. To this end, Abbasi-Yadkori & Szepesvári (2011) transformed the system into the following linear model:
@@
As noted by Tsiamis et al. (2023), finite-sample guarantees for system identification under heavy-tailed noise remain an open challenge. In light of Abbasi-Yadkori & Szepesvári (2011)’s application of linear bandit techniques to sub-Gaussian cases, we believe our method for HvtLB can be also benefit adaptive control in heavy-tailed settings. However, MOM methods are infeasible as the feature $z_{t}$, which includes the evolving state $x_{t}$, cannot be repeatedly sampled, and truncation-based methods rely on assumptions of bounded absolute moments for states, which may not hold. While adaptive Huber regression (Li & Sun, 2024; Huang et al., 2023) offers promising theoretical guarantees for heavy-tailed linear system identification, its computational inefficiency makes it unsuitable for adaptive control. In contrast, our proposed Hvt-UCB algorithm provides one-pass updates and has the potential to deliver finite-sample guarantees for adaptive control under heavy-tailed noise.
@@ 7 Conclusion
In this paper, we investigate the problem of heavy-tailed linear bandits (HvtLB). We highlight the advantages of Huber loss-based methods over truncation and median-of-means strategies, while also identifying the inefficiencies in previous Huber loss-based methods for HvtLB. To address these issues, we propose a Huber loss-based one-pass algorithm Hvt-UCB, based on Online Mirror Descent (OMD), a well-known regret minimization framework in online learning, which we adapt here for use as a statistical estimator. Hvt-UCB achieves the optimal and instance-dependent regret bound $\widetilde{\mathcal{O}}\big{(}dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\sqrt{\sum_{t=1}^{T}\nu_{t}^{2}}+dT^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\big{)}$ while only requiring $\mathcal{O}(1)$ per-round computational cost. Furthermore, our proposed method enjoys the potential to be extended to more broad online decision-making problems, including online linear MDPs and online adaptive control, thereby broadening its applicability. The key contribution is our adaptation of the OMD framework to stochastic linear bandits, effectively addressing the challenges introduced by heavy-tailed noise and the structure of the Huber loss.
Both our work and Huang et al. (2023) rely on prior knowledge of the moment bound $\nu_{t}$ to achieve instance-dependent guarantees. Recent progress in heavy-tailed multi-armed bandits has removed this requirement in different settings: adversarial (Huang et al., 2022), stochastic (Genalti et al., 2024), and best-of-both-worlds (Chen et al., 2025). Whether similar techniques can be applied to heavy-tailed linear bandits to relax this assumption remains an open question.
Acknowledgements
This research was supported by National Science and Technology Major Project (2022ZD0114800), NSFC (U23A20382), Postgraduate Research & Practice Innovation Program of Jiangsu Province (KYCX25_0327). The authors also thank Bo Xue for helpful discussions.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Properties of Huber Loss
Property 1 (Property 1 of Huang et al. (2023)).
@@ B.2 Proof of Lemma 3
$\begin{split}\frac{1}{2}\left\|\widehat{\theta}_{t}-\theta_{*}\right\|^{2}_{\frac{X_{t}X_{t}^{\top}}{\sigma_{t}^{2}}}\leq\left\langle\nabla\ell_{t}(\widehat{\theta}_{t}),\widehat{\theta}_{t+1}-\theta_{*}\right\rangle+\left\langle\nabla\ell_{t}(\widehat{\theta}_{t}),\widehat{\theta}_{t}-\widehat{\theta}_{t+1}\right\rangle+\left\langle\nabla\widetilde{\ell}_{t}(\widehat{\theta}_{t})-\nabla\ell_{t}(\widehat{\theta}_{t}),\widehat{\theta}_{t}-\theta_{*}\right\rangle,\end{split}$ (21)
where the first term can be bounded by the Bregman proximal inequality (Lemma 9), we have
@@ B.3 Proof of Lemma 4
By appling Lemma C.5 of (Huang et al., 2023), restated in Lemma 6, with $\tau_{t}=\tau_{0}\frac{\sqrt{1+w_{t}^{2}}}{w_{t}}t^{\frac{1-\varepsilon}{2(1+\varepsilon)}}$ and $b=\max_{t\in[T]}\frac{\nu_{t}}{\sigma_{t}}\leq 1$, we have with probability at least $1-\delta,\forall t\geq 1$,
@@ B.4 Proof of Lemma 5
where the last inequality comes from that $\max_{s\in[t]}\beta_{s-1}\leq\max_{s\in[t+1]}\beta_{s-1}$ Same as Eq. (C.12) of Huang et al. (2023), by setting $\tau_{0}$ as (17), with probability at least $1-\delta$, for all $t\geq 1$, we have
@@
where the last inequality comes from that $\operatorname{\mathds{1}}_{A_{s}}\leq 1$. Then, Lemma C.2 of Huang et al. (2023) (Self-normalized concentration) shows that by setting $\tau_{0}$ as (17) and $b=\max_{t\in[T]}\frac{\nu_{t}}{\sigma_{t}}\leq 1$, we have with probability at least $1-2\delta$, $\forall t\geq 1$,
@@ C.1 Proof of Theorem 1
where the last inequality comes from Lemma 7. Combining these two cases together, by choosing $\delta=\frac{1}{8T}$, $\lambda=d$, we have
@@ Appendix D Technical Lemmas
This section contains some technical lemmas that are used in the proofs.
@@ D.1 Concentrations
Theorem 2 (Self-normalized concentration for scalar (Abbasi-Yadkori et al., 2012, Lemma 7)).
@@
Lemma 6 (Lemma C.5 of Huang et al. (2023)).
@@ D.2 Potential Lemma
Note that in the last inequality, we use the fact that $\lVert U_{t-1}^{-\frac{1}{2}}X_{t}\rVert_{2}^{2}\leq c_{\max}$ and $1+x\geq\exp(x/2)$ holds for any $x\in[0,1]$. By taking advantage of the telescope structure, we have
@@
Lemma 8 (Lemma 5 of Zhao et al. (2020a)).
@@ D.3 Useful Lemma for OMD
Lemma 9 (Bregman proximal inequality (Chen & Teboulle, 1993, Lemma 3.2)).
@@
Appendix E Additional Experimental Results
In this section, we provide additional experimental results to validate the effectiveness of our algorithm. Appendix E.1 shows the experiment of different noise distributions, Appendix E.2 presents the experiment of varying $\nu_{t}$, and Appendix E.3 provides the experiment of varying arm set.
E.1 Experiment of Different Noise Distribution
(a) Pareto distribution
(b) Lomax distribution
(c) Fisk distribution
Figure 3: Experiment of Different Noise Distribution
Configurations.   We consider three classic heavy-tailed distributions: (a) Pareto distribution, (b) Lomax distribution, and (c) Fisk distribution. We implement all three distributions using Python’s scipy.stats. For the Pareto distribution, we set $b=1.5$; for the Lomax and Fisk distributions, we set $c=1.5$. We set $\varepsilon=1.49$, ensuring that the $(1+\varepsilon)$-moment exists. In the experiments, we use a time-varying noise scale, consistent with the setup in Figure 4.
Results.   We conduct 5 independent trials and averaged the results. Figure 3(a),  3(b) and  3(c) show cumulative regret under different noise distributions. In both settings, our algorithm performs similarly to Heavy-OFUL, demonstrating competitive performance. Note that the CRMM algorithm performs poorly across all three distributions, as it can only handle symmetric noise, while these distributions are non-symmetric. In summary, this experiment validates the effectiveness of our algorithm in handling various noise distributions.
E.2 Experiment of Varying $\nu_{t}$
(a) scale of $\nu_{t}$ is $1.5$
(b) scale of $\nu_{t}$ is $2$
(c) scale of $\nu_{t}$ is $2.5$
Figure 4: Experiment of Varying $\nu_{t}$.
Configurations.   In this experiment, we consider a Student’s t-distribution with a time-varying noise scale. Specifically, we fix the degrees of freedom as $df=1.7$ and set $\varepsilon=df-0.01$. At each round, noise is first sampled from Student $\text{Student t}(df)$ and then scaled by a factor $\alpha$, where $\log_{10}(\alpha)\sim\text{Unif}(0,scale)$, so that the central moments of $\epsilon_{t}$ vary across rounds. In this setting, existing algorithms can only rely on an upper bound of the noise variance, whereas our proposed HvtLB algorithm leverages the actual per-round variance for more accurate and adaptive scheduling. This setting follows the experimental setup proposed by Huang et al. (2023).
Results.   We test different scales of $\nu_{t}$, with scales of $1.5$, $2$ and $2.5$ in Figure 4(a),  4(b) and  4(c). In all environments, HvtLB and HeavyOFUL, which have variance-aware capabilities, outperform other algorithms significantly. In contrast, OFUL performs the worst because it lacks both variance-awareness and the ability to handle heavy-tailed noise.
E.3 Experiment of Varying Arm Set
(a) fixed arm set
(b) varying arm set 40/50
(c) varying arm set 30/50
Figure 5: Experiment of Varying Arm Set Case.
Configurations.   In this experiment, we create a total arm set $\mathcal{X}$ with 50 arms, which are pre-sampled before the experiment begins. To make the arm set varying, in each round, we randomly select a subset from $\mathcal{X}$ as the current arm set $\mathcal{X}_{t}$, from which the algorithm can choose arms. In this case, the regret is based on the best arm in each round, rather than the global optimal arm. We consider three scenarios: (i) the arm set $\mathcal{X}_{t}=\mathcal{X}$ with 50 arms, (ii) randomly selecting 40 arms in each round, and (iii) randomly selecting 30 arms in each round.
Results.   We conduct 5 independent trials and averaged the results. Figure 5(a),  5(b) and  5(c) show cumulative regret under different levels of arm set changes. In all settings, our algorithm performs the best, while HeavyOFUL and CRTM also achieve competitive results. However, when the arm set varies, the MOM-based algorithm fails. This is because MOM-based algorithms assume a fixed arm set and rely on resampling. When the arm set changes, resampling the same arms is not possible, so the algorithm selects the closest available arms instead. In summary, this experiment validates the effectiveness of our algorithm in scenarios where the arm set is time-varying.